# EEG-CLIP : Learning EEG representations from natural language descriptions

### v1

@@ 4.3.2 Projected embedding dimension
Additionally, we analyze EEG-CLIP model performance across varied hidden dimensionality sizes for the jointly learned EEG and text embeddings. Counter to typical representation learning trends, we find higher decoding accuracy with smaller shared embedding spaces between 32-128 dimensions rather than larger 256 or 512 sizes. Compressing the modalities into more compact unified vectors likely distills the patterns into their most essential latent characteristics necessary for generalization. Meanwhile, overparameterization could enable overfitting particular artifacts of the training distribution. Further, smaller spaces may inherently force more direct alignment between descriptive concepts and neurological patterns. By tuning towards optimal lower-complexity manifolds, EEG-CLIP better transfers key aspects in the aligned embeddings while filtering out unnecessary signals.
@@ 5 Discussion
The model achieves strong performance on standard classification tasks (balanced accuracies of 0.847 for pathology, 0.702 for gender, and 0.747 for age) when using an MLP classifier head. Most notably, EEG-CLIP demonstrates zero-shot classification capabilities, achieving 0.755 balanced accuracy on pathology detection using only natural language prompts without any task-specific training. In low-data regimes, our approach shows substantial benefits over models trained from scratch or pretrained on irrelevant tasks, suggesting efficient capture of generalizable features.
@@
Despite these promising results, several limitations warrant discussion. Our experiments were limited to the TUAB dataset, which may not capture the full diversity of EEG patterns and neurological conditions. Clinical reports in the dataset vary in quality, detail, and structure, potentially limiting the model’s ability to learn precise alignments. Additionally, our current approach treats entire EEG recordings and their corresponding reports as aligned pairs, whereas more fine-grained temporal alignment could further improve performance.
Future work could explore methods for aligning specific EEG segments with relevant sentences in clinical reports, scaling to larger and more diverse EEG datasets, incorporating additional data modalities, and developing methods to interpret the learned representations in terms of clinically meaningful EEG patterns. The development of foundation models for biosignals, analogous to large language models or vision transformers, represents an exciting direction for this research.
In conclusion, EEG-CLIP demonstrates the feasibility of contrastive learning between EEG signals and natural language descriptions for developing more general and transferable EEG representations. While this represents just an initial exploration, the approach opens up new possibilities for leveraging unstructured clinical text to enhance deep learning models for EEG analysis, potentially leading to more flexible, efficient, and powerful tools for neurological assessment and research.

### v2

@@ 4.3.2 Projected embedding dimension
Additionally, we analyzed EEG-CLIP model performance across varied hidden dimensionality sizes for the jointly learned EEG and text embeddings, as illustrated in Figure 8. Counter to typical representation learning trends, we found higher decoding accuracy with smaller shared embedding spaces between 32-128 dimensions rather than larger 256 or 512 sizes. A t-SNE visualization of the 64-dimensional embeddings in Figure 9 reveals clear clustering by pathological status, demonstrating effective semantic organization of the learned representation space.
This counterintuitive finding suggests that compressing both modalities into compact unified vectors distills patterns into their most essential characteristics necessary for generalization, while avoiding overfitting to training distribution artifacts that may occur in higher-dimensional spaces. The constrained dimensionality may also enforce more direct alignment between descriptive clinical concepts and underlying neurological patterns. These results indicate that EEG-CLIP benefits from lower-complexity manifolds that capture key cross-modal correspondences while filtering out extraneous signals.
4.4 Gradient Analysis for Model Interpretability
In order to provide insights into EEG-CLIP’s decision-making process, we performed gradient-based analysis to visualize which EEG regions contribute most strongly to embeddings aligned with specific textual concepts. We computed gradients of the cosine similarity between EEG embeddings and text embeddings for various prompts with regard to the EEG input that is forwarded through the EEG encoder. We computed those gradients with regard to the input in the frequency domain for easier analysis. Gradients were averaged across the validation set of TUAB.
Analysis of prompts containing specific frequency descriptions reveals characteristic gradient patterns. The prompt "Excessive beta activity" (Figure 10) produces elevated gradient responses in the beta frequency range (14-30 Hz).
For the more spatially specific prompt "20 Hz beta spindles in central regions" (Figure 11), gradients concentrate in the central electrode locations (C3, Cz, C4) with a peak at approximately 20 Hz, demonstrating the model’s ability to capture both spatial and frequency-specific features mentioned in the text.
The prompt "Left temporal sharp waves with a frequency of 6 Hz" (Figure 12) produces gradients with lateralized patterns, showing higher magnitudes in left temporal electrodes (T3, T5) with a prominent peak around 6 Hz, indicating successful spatial-frequency alignment with the textual description.
Figure 13 shows the gradient patterns when the model aligns EEG signals with "This is a normal recording" embeddings. The frequency gradient analysis reveals a positive peak around 8-9 Hz in the alpha range, indicating the model identifies healthy recordings with increased alpha band activity. This pattern suggests the model has learned to identify normal EEG patterns through alpha frequency features.
These gradient visualizations provide preliminary evidence that EEG-CLIP learns clinically relevant spatial-temporal patterns rather than purely relying on spurious correlations. The distinct gradient patterns between different prompts suggest the model captures meaningful neurophysiological differences. However, more sophisticated interpretability methods and validation with clinical experts would be needed to fully understand the clinical relevance of these learned representations.
@@ 5 Discussion
The model achieves strong performance on standard classification tasks (balanced accuracies of 0.847 for pathology, 0.702 for gender, and 0.747 for age) when using an MLP classifier head. Most notably, EEG-CLIP demonstrates zero-shot classification capabilities, achieving 0.755 balanced accuracy on pathology detection using only natural language prompts without any task-specific training. In low-data regimes, our approach shows substantial benefits over models trained from scratch or pretrained on alternative tasks, suggesting efficient capture of generalizable features.
@@
Despite these promising results, several limitations warrant discussion. A significant limitation is our primary reliance on TUAB for evaluation, which raises valid concerns about dataset-specific biases and overfitting. Clinical reports in the dataset vary in quality, detail, and structure, potentially limiting the model’s ability to learn precise alignments. Additionally, our current approach treats entire EEG recordings and their corresponding reports as aligned pairs, whereas more fine-grained temporal alignment between specific EEG segments and relevant report sections could improve performance, as demonstrated by recent Multiple Instance Learning frameworks (Gijsen & Ritter, 2024). While we implemented validation strategies including held-out evaluation sets, cross-validation for few-shot experiments, and evaluation across four diverse tasks, broader multi-site validation across different hospital systems, recording protocols, patient populations, and clinical practices would strengthen generalizability claims.
Future work could explore methods for aligning specific EEG segments with relevant sentences in clinical reports, scaling to larger and more diverse EEG datasets, incorporating additional data modalities, and developing methods to interpret the learned representations in terms of clinically meaningful EEG patterns. The field would benefit from standardized multi-site evaluation protocols for EEG-language models to better assess generalizability across diverse clinical settings.
In conclusion, EEG-CLIP demonstrates the feasibility of contrastive learning between EEG signals and natural language descriptions for developing more general and transferable EEG representations. While this work represents an initial exploration, the approach opens up new possibilities for leveraging unstructured clinical text to enhance deep learning models for EEG analysis, potentially leading to more flexible and data-efficient tools for neurological assessment and research.